OpenStudy (anonymous):

log(4)+log(x)=log(5)-log(x)

OpenStudy (anonymous):

use log(a*b) = log(a) + log(b) on the left. use log(a/b) = log(a) - log(b) on the right. what equation are you left with?

OpenStudy (anonymous):

log(4x)=log (5/x)

OpenStudy (anonymous):

since they're logs of the same base, you can just equate 4x = 5/x this is what you need to solve now..

OpenStudy (anonymous):

got it thanks...

OpenStudy (anonymous):

yw :)